At the center of a sphere of radius 0.5m a point charge of 2 nC is placed. What electric flux is passing through the sphere?


1
Expert's answer
2022-05-14T16:44:04-0400

The Gauss' law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

"\phi=\frac{2*10^{-9}}{8.85*10^{-12}}=226\:\rm N\cdot m^2/C"